### v3

# Optimization in quantum metrology and entanglement theory using semidefinite programming

Abstract
We discuss efficient methods to optimize the metrological performance over local Hamiltonians in a bipartite quantum system. For a given quantum state, our methods find the best local Hamiltonian for which the state outperforms separable states the most from the point of view of quantum metrology. We show that this problem can be reduced to maximize the quantum Fisher information over a certain set of Hamiltonians. We present the quantum Fisher information in a bilinear form and maximize it by an iterative see-saw (ISS) method, in which each step is based on semidefinite programming. We also solve the problem with the method of moments that works very well for smaller systems. We consider a number of other problems in quantum information theory that can be solved in a similar manner. For instance, we determine the bound entangled quantum states that maximally violate the Computable Cross Norm-Realignment (CCNR) criterion. Our approach is one of the efficient methods that can be applied for an optimization of the unitary dynamics in quantum metrology, the other methods being, for example, machine learning, variational quantum circuits, or neural networks. The advantage of our method is the fast and robust convergence due to the simple mathematical structure of the approach.
I Introduction
Quantum entanglement is at the heart of quantum physics and several quantum information processing applications [1, 2, 3, 4]. It also plays a central role in quantum metrology. It is known that interparticle entanglement is needed to overcome the shot-noise limit of the precision in parameter estimation in a linear interferometer [5]. It has also been shown that the larger the achieved precision in parameter estimation with a multiparticle system, the larger the depth of entanglement the state must possess [6, 7]. However, not all entangled states are useful for metrology [8].
At this point, several important questions arise. How can we find the best quantum state for metrology for a given setup? Perhaps, historically one of the first methods was the iterative see-saw (ISS) method applied in the quantum Fisher information picture [9, 10, 11, 12], as well as in the Bayesian one [13, 14, 15]. There are also methods based on an optimization over the purifications [16, 17, 18]. It is also possible to use variational techniques for efficiently searching in the high-dimensional space of quantum states, which was also applied for a similar task [19, 20]. The variational search methods can even be tailored for particular physical systems, such as tweezer arrays as programmable quantum sensors [21].
Another relevant question is how we could find the optimal Hamiltonian for a given quantum state. In the qubit case, it has been studied how to optimize the metrological performance for local Hamiltonians that are all Pauli spin matrices rotated by some unitary [8]. In this case, the best metrological performance achievable by separable states is a constant independent of the Hamiltonian. Any state that has a better metrological performance than that must be entangled.
The problem of the metrological performance of a bipartite state of qudits with a dimension larger than two is more complicated. In this case, the local Hamiltonians cannot be transformed into each other by local unitaries, and the best metrological performance achievable by separable states depends on the Hamiltonian. Thus, we introduce the metrological gain as a general way to characterize the quantum advantage of a system with arbitrary dimensions. It is defined as the ratio of the quantum Fisher information and the maximum of the quantum Fisher information over separable states maximized over local Hamiltonians [22]. The metrological gain is convex in the state, and it cannot decrease if additional copies or ancilla systems are added [22]. The optimization over the local Hamiltonian can also be carried out via an ISS method [22, 12, 23]. Thus, our method uses an ISS algorithm for optimizing the unitary dynamics rather than for optimizing the state. Note that instead of using the metrological gain, one can also consider the difference of the quantum Fisher information and its maximum for separable states [24], or even consider the difference with respect to a bound that depends on the state [25, 26, 27, 28].
In this paper, we examine the optimization problem above in a great depth. We elaborate the point that local Hamiltonians with meaningful constraints form a convex set. Thus, we have to maximize the metrological gain over a convex set of local Hamiltonians. We find several algorithms to maximize the metrological gain. The first method is based on maximizing a quadratic expression using a see-saw over a bilinear expression, and its computational costs are minimal. We present other algorithms that need quadratic programming with quadratic constraints and relaxations. Such methods work for small systems. We test the methods and compare them to each other. We list a number of similar problems that can be solved in an analogous way, such as maximization of norms over a convex set.
Our work is also related to recent efforts to solve various problems of numerical optimization in quantum metrology, including the combination of the tasks mentioned above, such as using an ISS for optimizing the quantum state and the unitary dynamics [22]. An ISS-based method has been used to optimize over a general dynamics using its Choi-Jamiolkowski representation [29]. An ISS-based method has also been used to optimize over adaptive strategies, when coherently probing several independent quantum channels [30]. An optimization over the probe state and the dynamics has been realized experimentally in a programmable quantum sensor realized in trapped cold ions [31]. For the identification of the optimal quantum metrological protocol in large systems, matrix product operators have been used [11]. Optimization methods, including ISS, has been considered in estimating spectral density of a stochastic signal field [32]. Finally, numerical optimization in quantum metrology in various complex scenarios has been considered including multiparameter estimation [33, 34, 35, 36, 37, 38], and there are even methods that are based on semidefinite programming [39].
Our work has a relation to resource theories, in which the quantum Fisher information can be used to identify quantum resources [24, 40]. Indeed, quantum Fisher information can be ideal to characterize quantum systems realized in Noisy Intermediate-Scale Quantum applications [41]. For a resource theory, we need to identify a convex set of free quantum states and a set of free operations. In our case, the free states are the separable states and the free operations include the local unitary operations. We can use the robustness of metrological usefulness defined in Ref. [10] to quantify our resource. It asks the question, how much noise can be added to the state such that it is not useful anymore metrologically. In this paper, we will examine another measure, the metrological gain mentioned above.
Finally, our work is relevant for the field that attempts to find a quantum advantage in experiments [42, 43, 44, 45]. The bound for quantum Fisher information for separable states has already been used to detect metrologically useful multipartite entanglement in several experiments [46, 47, 48].
Our paper is organized as follows. In Sec. II, we review the basic notions of entanglement theory and quantum metrology. In Sec. III we discuss, how our optimization problem can be viewed as maximization over a convex set of Hamiltonians. We review the notion of metrological gain. In Sec. IV, we present various methods to optimize the quantum Fisher information over a local Hamiltonian. In Sec. V, we consider optimization problems in quantum information that can be solved in a similar way. For instance, we determine the bound entangled quantum states that maximally violate the Computable Cross Norm-Realignment (CCNR) criterion for entanglement detection [49, 50].
II Background
In this section, we review entanglement theory and quantum metrology.
II.1 Entanglement theory
In this section, we summarize basic notions of entanglement theory [1, 2]. A bipartite quantum state is called separable if it can be written as mixture of product states as [51]
where are probabilities. If a quantum state cannot be decomposed as in Eq. (1), then the state is entangled.
Deciding whether a quantum state is entangled is a hard task in general. However, there are some necessary conditions for separability, that are easy to test. If these conditions are violated then the state is entangled.
One of the most important conditions of this type is the condition based on the positivity of the partial transpose (PPT). For a bipartite density matrix given as
the partial transpose according to first subsystem is defined by exchanging subscripts and as
It has been shown that for separable quantum states [52, 53]
holds. Thus, if has a negative eigenvalue then the quantum state is entangled. For and systems, the PPT condition detects all entangled states [53]. For systems of size and larger, there are PPT entangled states [54, 55]. They are also quantum states that are called bound entangled since their entanglement cannot be distilled to singlet states with local operations and classical communication (LOCC). Some of the bound entangled states are detected as entangled by the CCNR criterion [49, 50].
II.2 Quantum metrology
In this section, we summarize basic notions of quantum metrology such as the quantum Fisher information and the Cramér-Rao bound. We will also discuss the multi-parameter case.
The fundamental task of quantum metrology is the following. We have a probe state or an initial state that evolves according to the dynamics
where the unitary is given as
Here is the Hamiltonian of the evolution and is a parameter. Then, the metrological precision in estimating is bounded by the Cramér-Rao bound [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66]
where is the number of independent repetitions, and is the quantum Fisher information, defined as [56, 57, 58, 59, 60]
Here, we consider a density matrix with the following eigendecomposition
the constant coefficients depend only on the eigenvalues of the probe state and are given as
whenever the denominator is not zero, and the matrix element of the Hamiltonian in the basis of the eigenvectors of the density matrix is given as
The quantum Fisher information in Eq. (8) can be written in a different form that will be useful in our calculations as
where the superscripts "r" and "i" denote the real and imaginary parts, respectively.
After defining the quantum Fisher information, it is important to stress its basic properties, namely, that it is convex in the state, and it is bounded from above by the variance as
where for pure states there is an equality in Eq. (13).
Moreover, there is an important inequality for the quantum Fisher information with the error propagation formula [67, 68, 69]
which gives us, essentially, the precision of the parameter estimation, provided that we measure in the eigenbasis of the observable It can be shown that the uncertainty is the smallest, and thus the bound in Eq. (14) is the best if we choose to be the so-called symmetric logarithmic derivative
which fulfills
where is the anticommutator. In this case, the inequality in Eq. (14) is saturated. For pure states the symmetric logarithmic derivative is given as [64]
So far we discussed the case of estimating a single parameter. Let us consider now the estimation of several parameters when the unitary dynamics is given by
where is the Hamiltonian corresponding to the parameter The Cramér-Rao bound can be generalized to this case as
where the inequality in Eq. (19) means that the left-hand side is a positive semidefinite matrix, is now the covariance matrix with elements
and is the Fisher matrix, and its elements are given as
Here, the quantum Fisher information with two operators is given as
where denotes elementwise or Hadamard product. After straightforward algebra, the following equivalent formulation can be obtained
If has low rank, then, based on Eq. (10), the matrix is sparse, which can greatly simplify the calculation of Eq. (23). In contrast to the single parameter case, the generalized Cramér-Rao bound in Eq. (19) cannot be always saturated even in the limit of many repetitions.
III Basic notions for optimizing the metrological performance
When maximizing the quantum metrological performance over a Hamiltonian, we have to consider two important issues.
First, not all Hamiltonians are easy to realize. In particular, local Hamiltonians are much easier to realize than Hamiltonians with an interaction term, and such Hamiltonians appear most typically in quantum metrology. Thus, we will restrict the optimization for such Hamiltonians given as
where are single-subsystem operators. In the rest of the paper, will denote such local Hamiltonians. Moreover, we will use to denote the set of local Hamiltonians. With this, the considered metrological process is shown schematically in Fig. 1.
Second, we have to note that for the quantum Fisher information
holds, where is some constant. Thus, it is easy to improve the metrological performance of any Hamiltonian by multiplying it by a constant factor. In order to obtain a meaningful optimization problem, we have to consider a formulation of such a problem where such a "trick" cannot be used to increase the quantum Fisher information to infinity. One option could be to divide the quantum Fisher information with the square of some norm or seminorm of the Hamiltonian, for which
In principle, several norms for the normalization, for instance, the Hilbert-Schmidt norm
In this paper, we argue that it is advantageous to compare the metrological performance of a state to the best performance of separable states. Therefore, we will divide the quantum Fisher information of the state by the maximum quantum Fisher information achievable by separable states given as for some as [70, 10]
where and denotes the maximal and minimal eigenvalues, respectively.
III.1 Seminorm for local Hamiltonians
Let us now present the details of our arguments. We will examine some properties of the maximum of the quantum Fisher information for separable states given in Eq. (28). Let us consider first the quantity
[…]
With Eq. (29), the maximum of the quantum Fisher information for separable states, can be expressed as
Next, let us prove some important properties of
Observation 1. We know that is a seminorm. While this is shown in Ref. [71], for completeness and since the steps of the proof will be needed later, we present a brief argument to prove it.
Proof. We need the well-known fact that for Hermitian matrices and
hold, following from Weyl’s inequality [73]. Based on these, we will now prove that has all the properties of a seminorm [74, 75], examining the three properties one by one.
(ii) Absolute homogeneity holds since
for all and
(iii) Non-negativity holds since holds for all If is a zero matrix then However, does not imply that is a zero matrix, thus is not a norm, only a seminorm [74, 75].
Next, let us prove some important properties of the maximal quantum Fisher information for separable states.
Observation 2. The expression with the square root of the maximum for separable states
is a seminorm for
Proof. We will now follows steps similar to those in the proof of Observation 1, showing that the three properties of seminorms are satisfied.
(i) Let us consider a local Hamiltonian that is the sum of two local Hamiltonians
where the two local Hamiltonians are given as
Then, the following series of inequalities prove the subadditivity property
The first inequality in Eq. (38) is due to the subadditivity of and the monotonicity of in and The second inequality is based on the relation
which can be proved by straightforward algebra.
(ii) Absolute homogeneity,
for any real follows directly from the expression for with in Eq. (31), using the absolute homogeneity property of given in Eq (34)
(iii) Non-negativity can be proven using the expression for with in Eq. (31) and the non-negativity of .
Note that holds if and only if where is a real constant. Thus, in a certain sense, plays a role of distance between local Hamiltonians and such that the distance is zero if is the identity times some constant. In this case, the two matrices are not identical, but from the point of view of quantum metrology they act the same way.
Let us see some consequences of the observation. As any seminorm, is convex in the local Hamiltonian. Due to that, is also convex.
III.2 Metrological gain
Based on these ideas, we define the metrological gain for a given local Hamiltonian and for a given quantum state as [22]
In Eq. (41), we normalize the quantum Fisher information with a square of a seminorm of local Hamiltonians. Then, we would like to obtain the gain based on an optimization over all local Hamiltonians as
Maximizing over is difficult since both the numerator and denominator depend on Let us consider a subset of Hamiltonians of the type Eq. (24) such that
for Then, the maximum of the quantum Fisher information for separable states is
We can define the metrological gain for Hamiltonians that fulfill Eq. (43) as
where we call the set of local Hamiltonians satisfying Eq. (43). Finally, the maximal gain over all possible local Hamiltonians is given as
Note that is convex in [22].
Based on Eq. (45), we can see that computing is the same as maximizing the quantum Fisher information for the given set of Hamiltonians. We can do that by requiring that
where and is some constant. This way we make sure that
for Since we maximize a convex function over a convex set, the optimum is taken on the boundary of the set where the eigenvalues of are In this case,
Note that the constraints in Eq. (47) determine a convex set of Hamiltonians. That is, if two Hamiltonians, and , are of the form Eq. (24) with the constraints given in Eq. (47), then their convex combination, i.e.,
with is also of that form, as shown in Fig. 2. This can be seen from Eq. (32).
Thus, we need to deal with convex sets, similarly, as convex sets appear in the separability problem of quantum states [51, 1]. In Appendix A, in order to understand better the role of convexity in this case, we define a subset of separable quantum states similar to the local Hamiltonians we consider in this paper, and examine its properties.
In summary, we argue that the metrological gain, defined by Eq. (41), is a natural way to characterize the metrological performance of quantum states as it identifies states that perform better metrologically than all the separable states. Moreover, it considers local Hamiltonians, which are probably the least demanding to realize experimentally. Finally, we also showed that the maximum of the quantum Fisher information for separable states is an important quantity. Its square root is a seminorm over local Hamiltonians. The maximum of the quantum Fisher information for separable states is used in the definition of the metrological gain to normalize the quantum Fisher information.
We remark that there are other quantities to characterize the metrological performance. One can consider difference of the quantum Fisher information and the its maximum for separable states, and include the possibility of non-unitary dynamics from the point of view of resource theories. This way, all entangled states turn out to be a useful resource [24]. One can also consider the difference the quantum Fisher information and a bound that depends on the state, where the difference is positive only for entangled states [25, 26, 27, 28].
IV Method for maximizing the quantum Fisher information over local Hamiltonians
In the previous section, we have shown that maximizing the metrological gain is essentially reduced to maximizing the quantum Fisher information with some constraints on the local Hamiltonian given in Eq. (47), that is, computing
where we remember that is the set of local Hamiltonians fulfilling Eq. (43). Thus, in this section we will present methods that can maximize the quantum Fisher information with such constraints.
IV.1 See-saw optimization based on a bilinear form
In Ref. [22], a method was presented to maximize the quantum Fisher information over local Hamiltonians for a given quantum state. It is an iterative method, that needs to compute the symmetric logarithmic derivative Eq. (15) at each step and is based on quantum metrological considerations. In this section, we will present a simpler formulation based on general ideas in optimization theory. Later, this will also make it possible to find provably the global maximum of the quantum Fisher information.
The quantum Fisher information given in Eq. (12) is a weighted sum of squares of the real and imaginary parts of the elements of the Hamiltonian. It can further be written as
where contains the ’s and contains the ’s and the ’s. The matrix is a positive semidefinite diagonal matrix. Minimizing such an expression over is a simple problem. On the other hand, maximizing it is difficult. A usual way of maximizing such an expression is as follows [76, 77, 78]. We replace the optimization over by a simultaneous optimization of the vectors and of a bilinear form as
This can be a basis for a see-saw-type optimization.
Based on this, we can write the following.
Observation 3. The quantum Fisher information can be maximized with an iterative see-saw (ISS) method. We need to use the formula maximizing over local Hamiltonians
Based on Eq. (53), we can carry out the following algorithm. Initially, we choose randomly. Then, we maximize over the matrix while keeping fixed. Such an optimization can be carried out, for instance, by semidefinite programming [79] or by simple matrix algebraic calculations, see Appendix B. Then, we maximize over while keeping fixed. Then, we optimize again over , etc. This process is illustrated in Fig. 3. Note that we do not have to verify that the random initial Hamiltonian fulfills the constraints Eq. (47), since a random initial Hamiltonian and for any as initial Hamiltonian will lead to the same optimal (An alternative formulation of the optimization problem is in Appendix C. With similar methods, we also analyze the optimization over a quantum state rather than over the Hamiltonian Appendix D.)
Based on Eq. (23), we can see that Eq. (53) can be expressed as
This gives a meaningful physical interpretation of the see-saw method.
Since is convex in and if we maximize it over a convex set of Hamiltonians, then it will take the maximum at the boundary of the set. This optimization is a hard task, as we have discussed. We have to start from several random initial conditions, and only some of them will find a solution.
| Quantum state | Dim. | See-saw | Level | Level | See-saw | Level | Level |
| from Eq. (55) | 2 | 13.6421 | 13.6421 | 13.6421 | 10.3338 | 10.3338 | 10.3338 |
| 9 | 13.8828 | 14.0568 | - | 13.4330 | 13.6009 | - | |
| Horodecki, [55] | 3 | 4.8859 | 5.1280 | 4.8859 | 4.8020 | 4.9910 | 4.8020 |
| UPB [80, 81] | 3 | 5.6913 | 6.1667 | 5.6913 | 5.6913 | 6.1667 | 5.6913 |
| Met. useful PPT [10] | 4 | 9.3726 | 9.3726 | - | 9.3726 | 9.3726 | - |
| Private PPT [82, 12] | 6 | 10.1436 | 10.1436 | - | 10.1436 | 10.1436 | - |
We tested our method, together with the method of Ref. [22]. We used MATLAB [83] and semidefinite programming using MOSEK [84] and the YALMIP package [85]. We also used the QUBIT4MATLAB package [86, 87].
We tested the method for a number of quantum states. We considered the isotropic state of two -dimensional qudits defined as
where is the maximally entangled state and is the fraction of the white noise. In particular, we considered and . For such states, the optimal Hamiltonian is known analytically [22]. We also considered two copies of the state and given as where the bipartite system consists of the parties and We also considered some bound entangled states that have been examined in the literature, and are often used to test various algorithms. Such states have a nontrivial structure, are highly mixed, but can be metrologically useful. We looked at the state based on an Unextendible Product Basis (UPB) [80, 81], the Horodecki state for [55], the metrologically useful bound entangled state in Ref. [10], and the private state given in Ref. [82, 12]. The results are shown in Table 1. The success probability of the method of Ref. [22] is similar, however, the formulation used in this paper makes it possible to design methods that provably find the optimum 111We started from a random initial Hermitian Hamiltonian, where the real and imaginary parts of the elements had a normal distribution with a zero mean and a unit variance, and they were independent from each other, apart from the requirement of the Hermitianity..
IV.2 Quadratic programming with quadratic constraints
In the previous section, we considered methods that aim to maximize the quantum Fisher information over the Hamiltonian numerically. In practice, they are very efficient. Still, it is not certain that they find the global optimum and they might find a smaller value. In this section, we present methods that always find the global optimum. However, they work only for small systems. We also present simple approximations of such methods called, relaxation. They find either the maximum or a value larger than that. The method given in Sec. IV.1 looking for the maximum with the see-saw and the methods discussed in this section can be used together. The gap between the two results give information about the location of the maximum. If the two methods find the same value, then it is the global maximum.
The basic idea of the method of moments can be understood as follows. According to the Shor relaxation [89, *[][.OriginallypublishedinSov.J.Comput.Syst.Sci.25, 1-11(1987).]Shor1987Quadratic], the maximization of Eq. (51) is replaced by the maximization of
where the Hermitian is constrained as
Note that Eq. (57) can easily be coded in semidefinite programming. We can also add further and further conditions on that lead to lower and lower values for the maximum. Similar relaxation techniques have been used in entanglement theory to detect entanglement, non-locality [91], to obtain the maximum of an operator for separable states [92] or even to determine the ground state energy in spin one-half Heisenberg models [93].
Observation 4. The maximization of quantum Fisher information can be carried out with the method of moments. In this case, the conditions of Eq. (47) are replaced by
for , which makes sure that all eigenvalues of are The results of the first level relation can be improved if we add the constraints
for where where this way we set the number of ’s to a given value in the diagonal of . In quadratic programming, we can also use the constraint
which is equivalent to Eq. (58) for all values. Thus, we do not need to carry out calculations for a range of values. A somewhat weaker constraint, which is computationally simpler is
We tested our methods for the same states we considered in Sec. IV.1. For the example, on a laptop computer, we could carry out calculations at level and Even level gave the exact result. For the examples, we could carry out calculations at level and and level gave the correct result. For the and examples, the level calculation gave already the correct result. The results are shown in Table 1.
V Other problems in which sums of quadratic expressions must be maximized over a convex set
In this section, we consider optimization problems in quantum information science that can be formulated similarly to the maximization of the quantum Fisher information.
V.1 Optimization over various subsystems alternatingly
In quantum information, see-saw methods have already been used [95, 96, 97, 98, 99]. Problems related to entanglement in multipartite systems seem to be ideal for such methods. For instance, a typical problem is maximizing an operator expectation value for a product state
Then, we can maximize alternatingly over and over .
V.2 Maximization of the Wigner-Yanase skew information
The Wigner-Yanase skew information is defined as [100]
Here, we optimize over It is clear how to maximize Eq. (63) with a see-saw method.
Maximizing the Wigner-Yanase skew information can be used to lower bound the quantum Fisher information based on
The results are shown in Table 1. From a comparison with the results for the quantum Fisher information, we can see that for the last three of the five examples, there is an equality in Eq. (65). This has been known for the last two examples [10, 12]. For these states, we also have an equality in Eq. (13) with the optimal Hamiltonian, that is, the quantum Fisher information equals four times the variance. Surprisingly, for the fourth example, the UPB state, we do not have an equality in Eq. (13), which we can check with direct calculation.
V.3 Searching for the eigenstate with maximal/minimal eigenvalue
In this section, we discuss how to look for the largest eigenvalues of a matrix with a see-saw method similar to the one discussed in Sec. IV.1. Interestingly, the see-saw is guaranteed to converge in this case.
We need the relation that converts the maximization over a quadratic form into a maximization over a bilinear expression as
where is a Hermitian matrix and the right-hand side of Eq. (66) can immediately be used to define a see-saw algorithm. We shall show that such an algorithm is equivalent to the von Mises method (power iteration) for finding the maximal eigenvalue of a positive self-adjoint operator.
Before presenting the see-saw, let us consider the Cauchy-Schwarz inequality
where we took into account that . The inequality is saturated for
Now, the see-saw procedure is as follows:
(i) Choose some random unit vector.
(ii) Maximize over and alternatingly. First, is chosen based on Eq. (68). Then, in the alternate step, we use
We can now recognize the steps of the power iteration, which is known to converge as long as the starting vector ( in our case) overlaps with the eigenvector corresponding to the maximal eigenvalue of .
V.4 Maximization of the Hilbert-Schmidt norm
In this section, we look at the problem of maximizing a matrix norm with a method similar to the one discussed in Sec. IV.1. Let us consider the Hilbert-Schmidt norm given in Eq. (27), for which the maximization can be rewritten
where is some set of matrices. A straightforward see-saw optimization can be formulated based on Eq. (70). If is the set of density matrices then the maximization of the Hilbert-Schmidt norm is just the maximization of the purity or the minimization of the linear entropy
For instance, we can consider obtaining
where is a density matrix, are operators, and are constants. The constraints require that the expectation values of the operators equal the constants The optimization problem will lead to pure states if there are pure states satisfying the condition. It can be solved with the method of moments as well as with a see-saw iteration. The see-saw method shows a rapid convergence. In fact, it converges typically in a single step. After examining this simple case, we can have more complicated constraints on the quantum state. For instance, we can consider PPT states or states with a given negativity [10].
We add that in entanglement theory, the optimization over pure product states of bipartite and multipartite systems is an important task. This can be carried out if we require that the reduced density matrices are pure [92]. In this case, the purity of the reduced states appears among the constraints as
We add that optimization over pure states and rank constrained optimization is possible with other methods using semidefinite programming [101, 102].
V.5 Maximization of the trace norm
Another important norm in quantum physics, the trace norm is defined as
Observation 5. The trace norm of a Hermitian matrix given in Eq. (76) can be maximized with a see-saw method.
Proof. In order to obtain an optimization over a bilinear form, we use the dual relation between the trace norm and the spectral (operator) norm [103, 73], based on which
The maximization of over a set can then be written as
which can be calculated by a see-saw similar to the ones discussed before. We have to maximize alternatingly by and
In entanglement theory, the trace norm is used to define the CCNR criterion for entanglement detection [49, 50]. The criterion says that for every bipartite separable state i.e., states of the form given in Eq. (1), we have
where is the realigned matrix obtained by a certain permutation of the elements of If Eq. (79) is violated then the state is entangled. Clearly, the larger the left-hand side of Eq. (79), the larger the violation of Eq. (79).
The CCNR criterion can detect weakly entangled quantum states, called bound entangle states, that have a positive partial transpose. From this point of view, it is a very relevant question to search for PPT states which violate Eq. (79) the most. This can be done by maximizing over PPT states. We carried out a numerical optimization, using semidefinite programming. We present the maximum of for PPT states for various dimensions in Table 2. The state for the case could be determined analytically, see Appendix E.
| 1 | |||||
| 1 | |||||
| 1.1891 | |||||
| 1.2239 | |||||
| 1.5 | |||||
| 1.5 | |||||
| 1.5881 |
VI Conclusions
We discussed how to maximize the metrological gain over local Hamiltonians for a bipartite quantum system. We presented methods that exploit the fact that the the quantum Fisher information can be written as a bilinear expression of the elements of the Hamiltonian, if the Hamiltonian is given in the basis of the density matrix eigenstates. We showed an iterative see-saw (ISS) method that can maximize the quantum Fisher information for large systems. We showed another method based on quadratic programming. We tested the two approaches on concrete examples. We also considered several similar problems that can be solved based on such ideas, for instance, obtaining the maximum violation of the Computable Cross Norm-Realignment (CCNR) criterion for states with a positive partial transpose. Our work can straightforwardly be generalized to multiparticle systems.
VII Acknowledgements
We thank I. Apellaniz, R. Augusiak, R. Demkowicz-Dobrzański, O. Gühne, P. Horodecki, R. Horodecki, and J. Siewert for discussions. We thank J. Kołodyński, K. Macieszczak, and R. Demkowicz-Dobrzański for discussing the details of Ref. [9]. We acknowledge the support of the EU (COST Action CA23115, QuantERA MENTA, QuantERA QuSiED), the Spanish MCIU (Grant No. PCI2022-132947), the Basque Government (Grant No. IT1470-22), and the National Research, Development and Innovation Office NKFIH (Grant No. 2019-2.1.7-ERA-NET-2021-00036). We acknowledge the support of the Grant No. PID2021-126273NB-I00 funded by MCIN/AEI/10.13039/501100011033 and by "ERDF A way of making Europe". We thank the "Frontline" Research Excellence Programme of the NKFIH (Grant No. KKP133827). We thank the Project No. TKP2021-NVA-04, which has been implemented with the support provided by the Ministry of Innovation and Technology of Hungary from the National Research, Development and Innovation Fund, financed under the TKP2021-NVA funding scheme. We thank the Quantum Information National Laboratory of Hungary. G. T. thanks a Bessel Research Award of the Humboldt Foundation. We also acknowledge the support of the EU (QuantERA eDICT, CHIST-ERA MoDIC) and the National Research, Development and Innovation Office NKFIH (No. 2019-2.1.7-ERA-NET-2020-00003, No. 2023-1.2.1-ERA_NET-2023-00009, No. KH125096, and No. K145927).
Appendix A Special set of separable states similar to the set of local Hamiltonians
In this Appendix, we consider a set of separable states that are analogous to local Hamiltonians, namely states of the form
where and are the dimensions of the two Hilbert spaces.
First, for completeness, we explicitly show that the set of states of the form given in Eq. (80) is convex. Let
for , be two such states. Let us consider a density matrix being the mixture of and
where We shall show, that there exist , and such that can be written in the form given in Eq. (80). Let us consider the unnormalized density matrices
Let us also define their traces as
Then, with , can be formally written as in Eq. (80). We add that both and are density matrices: combinations of positive self-adjoint matrices with positive coefficients, thus positive, and of unit trace. What remains is to check that , which holds.
Let us now consider the recovery of and from a state of the form (80). As a first step, the reduced density matrices are
It is possible to solve Eqs. (85) for and as
However, if is not known, the recovery is not unique. Eqs. (86) yield a solution for any , for which and holds, constraining as
Note, that the bounds (87) also ensure that . Based on Eq. (87), a necessary condition for the existence of some allowed is
otherwise the state cannot be written in the form given in Eq. (80). It can also happen that Eq. (87) allows a single value. For instance, if one of the reduced density matrices or is not of full rank, only or , respectively, is possible.
Based on these, if a density matrix is given only, and one needs to know if it is of the form (80), one needs to choose a parameter fulfilling the bounds (87), calculate and , and afterwards verify, if formula (80) reproduces the density matrix given. If it works for one allowed value, it will work for the others as well, thus we need to try it only once. Thus, we can always decide whether a state is of the form Eq. (80), while it is not possible to decide efficiently whether a quantum state is separable or not [1, 2, 3, 4].
Straightforward algebra shows that for states of the form Eq. (81) and for traceless and Hermitian operators the relation
holds. Note that Eq. (89) does not mean that there are no correlations, since
is not necessarily zero. Let us now consider a full basis of traceless Hermitian matrices in the two subsystem
with the number of elements given as
for We assume that the basis matrices are pairwise orthogonal to each other
where and is the Kronecker symbol. Matrices and are the generators for and respectively. Then, any density matrix can be expressed as a linear combination as
for and Here, and are real coefficients that form the elements of the coherence vector of the system that describes the state of the system equivalently to the density matrix [104]. Due to Eq. (89), most of the elements of the coherence vector are zero since
holds for all Thus, the number of nonzero elements is at most
Appendix B See-saw algorithm with and without semidefinite programming
In this Appendix, we discuss how to implement a single step of the ISS method given in Sec. IV.1.
Based on Observation 3, we need to obtain
Here is the previous guess or the initial random guess. is of the form Eq. (24), where fulfill Eq. (47) for This can be solved by semidefinite programming.
Equation (96) can also be solved without semidefinite programming. The quantity to be maximized in Eq. (96) can be written as
where is given as
Let us write down the eigendecomposition of as
where is unitary and is diagonal. We define the diagonal matrix
where if and otherwise. With these the optimal Hamiltonians are given by
for From these, we can obtain the new and go to the next iteration step. Note that a similar scheme was presented in another context in Ref. [22].
Appendix C Alternative method to optimize quadratic functions
In this Appendix, we present an alternative of the method given in Sec. IV.1. We need the optimization over a constrained and an unconstrained quantity, while in the method of Sec. IV.1, we need to optimize over two constrained quantities.
Let us try to maximize the convex function
on a convex set. We could check the extreme point of the set, however, we look for a method that can easily be generalized for larger number of variables.
We introduce another function with an auxiliary variable that is easier to maximize. The linear function tangent to at the point is given as
where the quantity
is the value of the linear function at which equals times the Legendre transform of [105, 106, 107, 108, 109]. We can maximize given in Eq. (102) as
where is a convex set of real numbers, and is real. is plotted in Fig. 4 for
Then the maximization can be carried out as follows.
(i) Choose a random .
(ii) Maximize over for a fixed We have to use the condition for maximum
Hence, we have to set the auxiliary variable to
(iii) Maximize over for a fixed The maximum will be taken at an extreme point of After completing step (iii), go to step (ii).
The function is linear in concave in The Hessian matrix of the second order derivatives is
The eigenvalues of the Hessian are
One of the eigenvalues is positive for any Thus, is not concave in since the Hessian has positive eigenvalues. Thus, it is not guaranteed that we will find a unique maximum of a concave function.
Let us now generalize these ideas to several variables. Let us assume that and is a density matrix. Then, we can write that
where is defined as
Note the difference compared to the bilinear form in Eq. (53): The optimization over is unconstrained.
Appendix D Optimization over the state rather than the Hamiltonian
In this Appendix, we will analyze the method optimizing the quantum Fisher information over the quantum state, rather than over the Hamiltonian, described in Ref. [9, 110, 11]. We would like to examine whether it finds the global optimum. It turns out that the see-saw used for optimization is similar to that of Appendix C. Thus, we will apply an analysis similar to the one presented in Appendix C.
Let us review the method briefly. We need to maximize the error propagation formula
where the last equality stands, since operators for which maximize both and and for such operators Then, maximizing a fraction over such that appears both in the numerator and the denominator is a difficult task. Thus, a variational approach has been used to obtain the optimum as [9]
For the maximum [9]
holds. Then, for the optimum and substituting this into Eq. (LABEL:eq:fmin), we obtain the right-hand side of Eq. (112). If we carry out the optimization over and then we find that the optimum is taken when equals the symmetric logarithmic derivative and
In order to decrease the number of parameters we need to optimize, the optimization in Eq. (LABEL:eq:fmin) can be rewritten as
where takes the role of Note that the function in Eq. (115) is concave in and linear in
The expression in Eq. (115) can be used to define a two-step see-saw algorithm that will find better and better quantum states. First we choose a random state. Then, maximize the expression over One can show that the maximum is taken by setting to be the symmetric logarithmic derivative given in Eq. (15) [9]. Then, we optimize over Then, we go back to the step maximizing over and so on. The optimization seems to work very well, converging almost always in practice.
After reviewing the powerful method in Refs. [9, 110, 11], the question arises: do we always find the global maximum? For that we need that the expression maximized in Eq. (115) is strictly concave in Let us see a concrete example. Let us consider a single qubit with a density matrix
where the allowed values for the real parameter are given by
and is a subset of the interval. Clearly,
We will carry out an optimization of the quantum Fisher information over the convex set of states defined above. Let us now consider the Hamiltonian
Then, the symmetric logarithmic derivative, defined in Eq. (15), multiplied by is of the form
where is the auxiliary, real parameter. The expression maximized in Eq. (115) equals
which is maximized for a given if Then, we can eliminate and obtain
The Hessian matrix of the second order derivatives of the expression in Eq. (121) is
Since one of the two eigenvalues of the Hessian is negative, the other is positive, the expression in Eq. (121) is not concave in and In fact, Eq. (121) is concave in and linear in but not concave in The function given in Eq. (121) is plotted in Fig. 5, which makes it possible to understand the convexity properties of Eq. (121). Thus, the expression maximized in Eq. (115) is in general not concave in
Following the ideas of Appendix C and applying the Legendre transform [105, 106, 107, 108, 109], the equation of the tangent to the function given in Eq. (122) is
One can use an optimization over to obtain
Now, by substituting into Eq. (124) we arrive at Eq. (121). Following the ideas of Appendix C, we can also see that Eq. (115) is also based on linearizing at a given which leads to a tangent plane, and then maximizing the quantum Fisher information over that tangent plane.
Let us consider a maximization of Eq. (121) without further constraints such that and We start from a random which in our case means to start with a random Let us consider the case that Then, we will have Then, when optimizing over in the next step, we get Let us now consider the case that Then, we will have Then, when optimizing over we get Thus, we will find the correct maximum, independently from the initial state. The reason is that the Fisher information is invariant under the transformation
Observation 6. In the single-qubit case, the see-saw always reaches the global maximum for any initial state, if we optimize over the full set of physical quantum states.
Proof. Let us consider a single qubit in a general state
and a general single-qubit as
With these, the quantum Fisher information can be given as
The maximum is taken at and After eliminating we obtain
which takes the global maximum for all points for which and Based on these we see that starting from any random initial state, we always reach the global maximum, except for the case when when we obtain However, such states form a zero measure set, thus small additive noise with a random quantum state leads out of the set, and then the algorithm finds the global maximum.
In the more general case of larger systems, the quantum Fisher information is not the quadratic function of the density matrix elements anymore [111, 112]. However, even in this case, if is maximizing the quantum Fisher information then
also maximizes it, where Thus, even though the final state of the iteration depends on the initial state, the algorithm finds the global maximum. We tested the optimization problem for physical quantum states of various sizes (, , …, ). In this case, looking for the maximum of in Eq. (115) means identifying the eigenstate with the maximal eigenvalue of the operator [9]
For a pure state the symmetric logarithmic derivative for a given Hamiltonian is given in Eq. (17). Substituting it into , we obtain
where
If the initial state is an eigenvector of then i. e., a matrix with zero elements. Thus, the next state of the iteration obtained as the eigenvector of with the largest eigenvalue will be a random vector.
If the initial state is
then we have
Here, and denote the eigenvalues of while and are corresponding eigenvectors. The eigenvector of Eq. (134) with the largest eigenvalue is the state given in Eq. (133).
Let us see now the steady states of the see-saw. For such states,
holds, where is the maximal eigenvalue of given in Eq. (132). Thus, every steady state must fulfill the equation of the given Eq. (135) for some Let us consider first the case that has a non-generate spectrum. Let us look for the eigenvectors of the operator
where is a constant. For the solution is of the form where and are constants. For these states, For all all eigenstates of are solutions. The solutions for which are states of the type and states given in Eq. (133).
The case of we discussed in relation to Eq. (132). We find that states of the type given in Eq. (133) are stable points of the iteration. However, based on numerics, we find that a small perturbation outside of the subspace will move the iteration away from the state, unless we are in the state
which is stable under perturbations. Here, is a real number and and are the eigenvectors with the minimum and maximum eigenvalues of If has a degenerate maximal or minimal eigenvalue, then the steady states can also be obtained straightforwardly.
The steady states maximize
Thus, the mechanism of the see-saw resembles the power iteration discussed in Sec. V.3, however, now we maximize the variance of an operator. It has been shown that the iteration leads to states with a variance that does not decrease [9]. Together with our argument that the steady states all maximize the variance, we can see that the iteration always finds the global maximum. We verified this claim based on extensive numerical tests 222In principle, it is possible to find an initial state for which calculated by Eq. (15), equals the zero matrix. This is the case if is not in the range of the initial density matrix. However small random added noise can easily help, and the iteration will lead to the global maximum..
Next, we can consider the set of quantum states obtained via a physical map [9]. The first example is the states from the globally depolarizing map. We have to look for the eigenvector with the maximum eigenvalue in the following eigenequation [9]
where for computing , we used the symmetric logarithmic derivative of the noisy input state given as
The symmetric logarithmic derivative for is given as where is the symmetric logarithmic derivative for the noiseless pure state, and corresponds to no noise. Hence, we obtain
c. f. Eq. (131). The adjoint map is
Then, we arrive at
Hence, substituting with the symmetric logarithmic derivative given for pure states in Eq. (17), for we obtain
The part that contributes to finding the eigenvector is
c. f. Eq. (135). After a derivation similar to the previous one, we find again that maximizing and the same states maximize it as before. Thus, based on the same arguments, the algorithm always finds the global maximum.
Next, we present a case when the algorithm might not find the global optimum, while admittedly the example is not a very practical one, and is constructed to test the method. Let us consider a concrete iteration on a subset of physical quantum states of the type given in Eq. (116), and assume , which can be incorporated in the optimization as the condition
The inequality in Eq. (146) can be added as conditions to an optimization using semidefinite programming. Alternatively, we can also define a set of quantum states fulfilling Eq. (146) via the map
from the set of quantum states, where and are projectors to the eigenstates of The minimal and maximal value for respectively, are
Based on these, we can obtain the probabilities given with and as and
We start from a random that is, we start with a random Let us consider the case that Then, we will have Then, when optimizing over in the next step, we get Let us now consider the case that Then, we will have Then, when optimizing over we get Thus, depending on the sign of the random initial we will arrive at or at which is not necessarily the global optimum.
Considering a single qubit in a general state, conditions such as
where and are real numbers giving the bounds of the allowed region for , can result in reaching different local optima starting from different initial states, even if small random perturbations are added. The proof is similar to the proof of the one-dimensional case discussed before. For example if and the initial state is
then the optimization leads to and, based on Eq. (129), we have Thus, it does not reach the maximal value
The inequalities in Eq. (149) can be added as conditions to an optimization using semidefinite programming. Alternatively, we can also define the a set of quantum states via the physical map, e. g.,
where is the projector to the eigenstate of with eigenvalue and is the projector to the eigenstate of with eigenvalue For states coming from this map,
holds. If the optimization is carried out over the quantum states coming from the map then the maximum found will depend on the initial state.
More practical, relevant subsets of physical quantum states are PPT states in a bipartite system. Metrology with such states have been studied using semidefinite programming [10]. A family of bipartite PPT entangled states have been found numerically maximizing the quantum Fisher information [10]. Here, semi-definite programming was used to maximize over the set of PPT quantum states. Later, PPT states having the same quantum Fisher information have been found analytically, which is a strong indication that the method found the global maximum [12].
In multiparticle systems, one can consider states that are PPT with respect to all bipartitions. Such four-qubit states have been studied [10]. Such states form a very complicated set, and thus they are good candidates to test the method with a difficult task. We find that the maximum found in such systems might depend on the initial state. However, after starting from a couple of random initial values the iteration finds the global maximum.
One might think that the semidefinite solvers can cause the dependence on the initial states. It is possible to consider a subset of physical quantum states such that semidefinite solvers are not needed. Such a set is the set of states with a symmetric extension of a given order [114]. We considered the optimization over 2-symmetric extendible states and we also found a dependence on the initial state [115].
The optimization over PPT states can be reformulated with PPT-inducing channels acting on the set of quantum states [116], which are similar to entanglement breaking maps [117]. The set of the 2-symmetric extendible states can be obtained from the set of all quantum states with a physical map.
In summary, in most relevant multi-qudit problems, the method seems to converge extremely well, which is very likely due to the fact that the global maximum is taken by many states. It is also important that several auxiliary variables are added. When considering states passing through a physical map, the symmetry properties of the map determine how well the method converges. It would be interesting to prove that the approach always converges for several classes of relevant problems.
Appendix E Bound entangled state violating the CCNR criterion maximally
In this Appendix, we present the bound entangled state for which the violation of the CCNR criterion given in Eq. (79) is maximal. The state is
where the probabilities are
where the four Bell states are defined as
and the components for and are given as
For the state in Eq. (153), we have
Using the method of Sec. IV, we find that the state is not useful metrologically. We created a MATLAB routine that defines presented in this paper. It is part of the QUBIT4MATLAB package [86, 87]. The routine BES_CCNR4x4.m defines the state given in Eq. (153). We also included a routine that shows the usage of BES_CCNR4x4.m. It is called example_BES_CCNR4x4.m.

### v4

@@
-Optimization in quantum metrology and entanglement theory using semidefinite programming
+Iterative optimization in quantum metrology and entanglement theory using semidefinite programming
 Abstract
@@
 For a given quantum state, our methods find the best local Hamiltonian for which the state outperforms separable states the most from the point of view of quantum metrology.
-We show that this problem can be reduced to maximize the quantum Fisher information over a certain set of Hamiltonians.
+We show that this problem can be reduced to maximizing the quantum Fisher information over a certain set of Hamiltonians.
 We present the quantum Fisher information in a bilinear form and maximize it by an iterative see-saw (ISS) method, in which each step is based on semidefinite programming.
 We also solve the problem with the method of moments that works very well for smaller systems.
-We consider a number of other problems in quantum information theory that can be solved in a similar manner.
-For instance, we determine the bound entangled quantum states that maximally violate the Computable Cross Norm-Realignment (CCNR) criterion.
 Our approach is one of the efficient methods that can be applied for an optimization of the unitary dynamics in quantum metrology, the other methods being, for example, machine learning, variational quantum circuits, or neural networks.
 The advantage of our method is the fast and robust convergence due to the simple mathematical structure of the approach.
+We also consider a number of other problems in quantum information theory that can be solved in a similar manner.
+For instance, we determine the bound entangled quantum states that maximally violate the Computable Cross Norm-Realignment (CCNR) criterion.
 I Introduction
@@
 However, not all entangled states are useful for metrology [8].
+Even bound entangled states, which are considered very weakly entangled, can be more useful for metrology than some distillable entangled states [9, 10, 11].
 At this point, several important questions arise.
@@
 The metrological gain is convex in the state, and it cannot decrease if additional copies or ancilla systems are added [23].
-The optimization over the local Hamiltonian can also be carried out via an ISS method [22, 12, 23].
+In the multiparticle case, the metrological gain is related to multiparticle entanglement, as it lower bounds the entanglement depth [24].
+The optimization over the local Hamiltonians can also be carried out via an ISS method [23, 11, 24].
 Thus, our method uses an ISS algorithm for optimizing the unitary dynamics rather than for optimizing the state.
@@
 We test the methods and compare them to each other.
-We list a number of similar problems that can be solved in an analogous way, such as maximization of norms over a convex set.
+In the second half of the paper, we explore other important applications, where ISS methods can help with optimization.
+Namely, we list a number of similar problems that can be solved in an analogous way, such as maximization of norms over a convex set, maximizing the Wigner-Yanase skew information, looking for the extremal eigenvalue of Hermitian matrices, and also for looking for states that violate maximally the Computable Cross Norm-Realignment (CCNR) criterion [30, 31].
 Our work is also related to recent efforts to solve various problems of numerical optimization in quantum metrology, including the combination of the tasks mentioned above, such as using an ISS for optimizing the quantum state and the unitary dynamics [23].
-An ISS-based method has been used to optimize over a general dynamics using its Choi-Jamiolkowski representation [29].
-An ISS-based method has also been used to optimize over adaptive strategies, when coherently probing several independent quantum channels [30].
+An ISS-based method has been used for an optimization over a general dynamics using its Choi-Jamiołkowski representation [32].
+An ISS-based method has also been used for optimizing over adaptive strategies, when coherently probing several independent quantum channels [33].
 An optimization over the probe state and the dynamics has been realized experimentally in a programmable quantum sensor realized in trapped cold ions [34].
 For the identification of the optimal quantum metrological protocol in large systems, matrix product operators have been used [13].
-Optimization methods, including ISS, has been considered in estimating spectral density of a stochastic signal field [32].
-Finally, numerical optimization in quantum metrology in various complex scenarios has been considered including multiparameter estimation [33, 34, 35, 36, 37, 38], and there are even methods that are based on semidefinite programming [39].
+Optimization methods, including ISS, has been considered in estimating the spectral density of a stochastic signal field [35].
+Finally, numerical optimization in quantum metrology in various complex scenarios has been considered including multiparameter estimation [36, 37, 38, 39, 40, 41], and there are even methods that are based on semidefinite programmingx [42].
+Variational methods used for the optimization of the probe state have already been mentioned.
+General optimization methods have also been applied for the optimization of the dynamics and other complex tasks.
+Optimization of the controls has been carried out using neural networks and reinforcement learning [43].
+Machine learning approaches have been used in Bayesian parameter estimation [44].
+The advantageous aspects of ISS methods to general optimization methods are discussed in Ref.
+[45].
 Our work has a relation to resource theories, in which the quantum Fisher information can be used to identify quantum resources [25, 46].
@@
 In Sec.
-III we discuss, how our optimization problem can be viewed as maximization over a convex set of Hamiltonians.
+III, we discuss, how our optimization problem can be viewed as a maximization over a convex set of Hamiltonians.
 We review the notion of metrological gain.
@@
 V, we consider optimization problems in quantum information that can be solved in a similar way.
-For instance, we determine the bound entangled quantum states that maximally violate the Computable Cross Norm-Realignment (CCNR) criterion for entanglement detection [49, 50].
+For instance, we determine the bound entangled quantum states that maximally violate the CCNR criterion for entanglement detection [30, 31].
 II Background
@@
 II.1 Entanglement theory
-In this section, we summarize basic notions of entanglement theory [1, 2]. A bipartite quantum state is called separable if it can be written as mixture of product states as [51]
+In this section, we summarize basic notions of entanglement theory [1, 2]. A bipartite quantum state is called separable if it can be written as a mixture of product states as [55]
 where are probabilities. If a quantum state cannot be decomposed as in Eq. (1), then the state is entangled.
@@
 where the unitary is given as
-Here is the Hamiltonian of the evolution and is a parameter. Then, the metrological precision in estimating is bounded by the Cramér-Rao bound [56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66]
+Here is the Hamiltonian of the evolution, and is a parameter. Then, the metrological precision in estimating is bounded by the Cramér-Rao bound [60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70]
 where is the number of independent repetitions, and is the quantum Fisher information, defined as [60, 61, 62, 63, 64]
 Here, we consider a density matrix with the following eigendecomposition
-the constant coefficients depend only on the eigenvalues of the probe state and are given as
-whenever the denominator is not zero, and the matrix element of the Hamiltonian in the basis of the eigenvectors of the density matrix is given as
+the constant coefficients depend only on the eigenvalues of the probe state, and are given as
+whenever the denominator is not zero. The matrix elements of the Hamiltonian in the basis of the eigenvectors of the density matrix are given as
 The quantum Fisher information in Eq. (8) can be written in a different form that will be useful in our calculations as
-where the superscripts "r" and "i" denote the real and imaginary parts, respectively.
+where the superscripts r and i denote the real and imaginary parts, respectively.
 After defining the quantum Fisher information, it is important to stress its basic properties, namely, that it is convex in the state, and it is bounded from above by the variance as
@@
 First, not all Hamiltonians are easy to realize.
-In particular, local Hamiltonians are much easier to realize than Hamiltonians with an interaction term, and such Hamiltonians appear most typically in quantum metrology.
+In particular, local Hamiltonians are much easier to realize with a high precision than Hamiltonians with an interaction term.
+Local Hamiltonians appear most typically in quantum metrology forming the basis of linear interferometers.
 Thus, we will restrict the optimization for such Hamiltonians given as
@@
 Second, we have to note that for the quantum Fisher information
-holds, where is some constant.
+holds, where is a constant.
 Thus, it is easy to improve the metrological performance of any Hamiltonian by multiplying it by a constant factor.
@@
 In principle, several norms for the normalization, for instance, the Hilbert-Schmidt norm
+can be considered.
 In this paper, we argue that it is advantageous to compare the metrological performance of a state to the best performance of separable states.
 Therefore, we will divide the quantum Fisher information of the state by the maximum quantum Fisher information achievable by separable states given as for some as [74, 10]
-where and denotes the maximal and minimal eigenvalues, respectively.
+where and denote the maximal and minimal eigenvalues, respectively.
 III.1 Seminorm for local Hamiltonians
@@
 is a seminorm for
-Proof. We will now follows steps similar to those in the proof of Observation 1, showing that the three properties of seminorms are satisfied.
+Proof. We will now follow steps similar to those in the proof of Observation 1, showing that the three properties of seminorms are satisfied.
 (i) Let us consider a local Hamiltonian that is the sum of two local Hamiltonians
@@
 Then, we would like to obtain the gain based on an optimization over all local Hamiltonians as
+Note that is invariant under local unitaries thus
+where the transformed state is defined as
+Many entanglement measures, for instance, the Entanglement of Formation has the above property [1].
+We consider local Hamiltonians, as we have discussed at the beginning of Sec.
+III.
+While we do not study it in this work, it would also be interesting to consider non-local Hamiltonians for a similar definition.
+However, computing the analogous quantity for that case is difficult, since calculating the maximum of an operator expectation value for separable states is a hard task [1].
+Possibly, one can try to calculate the maximum for quantum states with a positive partial transpose (PPT) instead, which is an upper bound on the maximum for separable states, and it can be obtained via semidefinite programming (e.
+g., Refs.
+[80, 10]).
 Maximizing over is difficult since both the numerator and denominator depend on Let us consider a subset of Hamiltonians of the type Eq. (24) such that
@@
 The maximum of the quantum Fisher information for separable states is used in the definition of the metrological gain to normalize the quantum Fisher information.
+The metrological gain is also defined in the multiparticle case. It has been shown that, for the metrological gain the inequality [24]
+holds, where is the entanglement depth.
+It is assumed that the local Hamiltonians have identical maximal and minimal eigenvalues, which is fulfilled in typical many-particle systems [24].
+Thus, the advantage of using the metrological gain to characterize the performance of the setup is that it is directly related to multipartite entanglement.
+Moreover, maximal metrological gain is possible only with a maximal entanglement depth.
 We remark that there are other quantities to characterize the metrological performance.
-One can consider difference of the quantum Fisher information and the its maximum for separable states, and include the possibility of non-unitary dynamics from the point of view of resource theories.
+One can consider the difference of the quantum Fisher information and its maximum for separable states, and include the possibility of non-unitary dynamics from the point of view of resource theories.
+Considering the difference rather than the fraction makes the mathematical treatment easier.
 This way, all entangled states turn out to be a useful resource [25].
-One can also consider the difference the quantum Fisher information and a bound that depends on the state, where the difference is positive only for entangled states [25, 26, 27, 28].
+One can also consider the difference of the quantum Fisher information and a bound that depends on the state, where the difference is positive only for entangled states [26, 27, 28, 29].
+The advantage of this approach is that the quantity can be computed analytically.
 IV Method for maximizing the quantum Fisher information over local Hamiltonians
@@
 where contains the ’s and contains the ’s and the ’s.
-The matrix is a positive semidefinite diagonal matrix.
-Minimizing such an expression over is a simple problem.
-On the other hand, maximizing it is difficult.
+The matrix is a positive semidefinite symmetric matrix with real values.
+For the case of the quantum Fisher information, it is even a diagonal matrix.
+We will also consider other problems where it is a general positive semidefinite symmetric matrix.
+Minimizing such an expression over is a problem that can efficiently be solved.
+In general, minimizing a convex function under constraints is a task appearing often in quantum metrology.
+For instance, there are efficient ways to obtain the minimum of the quantum Fisher information over a given set of states [81, 82, 83].
+On the other hand, maximizing the expression given in Eq.
+(54) over is difficult.
 A usual way of maximizing such an expression is as follows [84, 85, 86].
@@
 This can be a basis for a see-saw-type optimization.
-Based on this, we can write the following.
+For completeness, let us prove Eq.
+(55) for the case of being a general positive semidefinite symmetric matrix.
+It is easy to see that the right-hand side is never smaller than the left-hand side.
+Now let us prove that the right-hand side is never larger than the left-hand side.
+Let us start from
+Hence, follows that
+where the equality is due to the fact that is symmetric. Then, at least one of the following two inequalities hold
+After including an optimization over and , the equality given in Eq. (55) follows.
+Based on the relation in Eq. (55), we can write the following.
 Observation 3. The quantum Fisher information can be maximized with an iterative see-saw (ISS) method. We need to use the formula maximizing over local Hamiltonians
@@
 3.
-Note that we do not have to verify that the random initial Hamiltonian fulfills the constraints Eq.
+Note that we do not have to verify that the random initial Hamiltonian fulfills the constraints given in Eq.
 (49), since a random initial Hamiltonian and for any as initial Hamiltonian will lead to the same optimal (An alternative formulation of the optimization problem is in Appendix C.
@@
 This optimization is a hard task, as we have discussed.
-We have to start from several random initial conditions, and only some of them will find a solution.
+We have to start from several random initial conditions, and only in some of the cases we will find a solution.
 | Quantum state | Dim. | See-saw | Level | Level | See-saw | Level | Level |
@@
 For such states, the optimal Hamiltonian is known analytically [23].
-We also considered two copies of the state and given as where the bipartite system consists of the parties and We also considered some bound entangled states that have been examined in the literature, and are often used to test various algorithms.
+We also considered two copies of the state with and given as where the bipartite system consists of the parties and We also considered some bound entangled states that have been examined in the literature, and are often used to test various algorithms.
 Such states have a nontrivial structure, are highly mixed, but can be metrologically useful.
@@
 The success probability of the method of Ref.
-[22] is similar, however, the formulation used in this paper makes it possible to design methods that provably find the optimum 111We started from a random initial Hermitian Hamiltonian, where the real and imaginary parts of the elements had a normal distribution with a zero mean and a unit variance, and they were independent from each other, apart from the requirement of the Hermitianity..
+[23] is similar, however, the formulation used in this paper makes it possible to design methods that provably find the optimum 111We started from a random initial Hermitian Hamiltonian, where the real and imaginary parts of the elements had a normal distribution with a zero mean and a unit variance, and they were independent from each other, apart from the requirement of the Hermiticity..
 IV.2 Quadratic programming with quadratic constraints
@@
 However, they work only for small systems.
-We also present simple approximations of such methods called, relaxation.
+We also present simple approximations of such methods called relaxation.
 They find either the maximum or a value larger than that.
@@
 The basic idea of the method of moments can be understood as follows.
-According to the Shor relaxation [89, *[][.OriginallypublishedinSov.J.Comput.Syst.Sci.25, 1-11(1987).]Shor1987Quadratic], the maximization of Eq.
+According to the Shor relaxation [97, *[][.OriginallypublishedinSov.J.Comput.Syst.Sci.25, 1-11(1987).]Shor1987Quadratic], the maximization of Eq.
 (54) is replaced by the maximization of
@@
 We can also add further and further conditions on that lead to lower and lower values for the maximum.
-Similar relaxation techniques have been used in entanglement theory to detect entanglement, non-locality [91], to obtain the maximum of an operator for separable states [92] or even to determine the ground state energy in spin one-half Heisenberg models [93].
+Similar relaxation techniques have been used in entanglement theory to detect entanglement, non-locality [99], to obtain the maximum of an operator for separable states [100], or even to determine the ground state energy in spin one-half Heisenberg models [101].
+For our problem, we used the method of moments [102].
 Observation 4. The maximization of quantum Fisher information can be carried out with the method of moments. In this case, the conditions of Eq. (49) are replaced by
@@
 for where where this way we set the number of ’s to a given value in the diagonal of . In quadratic programming, we can also use the constraint
-which is equivalent to Eq. (58) for all values. Thus, we do not need to carry out calculations for a range of values. A somewhat weaker constraint, which is computationally simpler is
+After simple considerations, we can find that Eq.
+(66) is equivalent to Eq.
+(64) for all values.
+Thus, we do not need to carry out calculations for a range of values.
+A somewhat weaker constraint, which is computationally simpler is
 We tested our methods for the same states we considered in Sec.
@@
 The results are shown in Table 1.
+A detailed explanation of the method via a concrete example is in Appendix E.
 V Other problems in which sums of quadratic expressions must be maximized over a convex set
@@
 Here, we optimize over It is clear how to maximize Eq. (69) with a see-saw method.
-Maximizing the Wigner-Yanase skew information can be used to lower bound the quantum Fisher information based on
+Here, instead of we use since the latter is Hermitian.
+The expectation value of a non-Hermitian matrix could be complex.
+Then, the method finds the maximum based on the derivation presented in Eqs.
+(56), (57), and (58).
+Maximizing the Wigner-Yanase skew information can be used to lower bound the quantum Fisher information based on [108]
 The results are shown in Table 1.
@@
 V.3 Searching for the eigenstate with maximal/minimal eigenvalue
-In this section, we discuss how to look for the largest eigenvalues of a matrix with a see-saw method similar to the one discussed in Sec.
+In this section, we discuss how to look for the largest eigenvalue of a positive semidefinite Hermitian matrix with a see-saw method similar to the one discussed in Sec.
 IV.1.
@@
 We need the relation that converts the maximization over a quadratic form into a maximization over a bilinear expression as
-where is a Hermitian matrix and the right-hand side of Eq.
-(66) can immediately be used to define a see-saw algorithm.
-We shall show that such an algorithm is equivalent to the von Mises method (power iteration) for finding the maximal eigenvalue of a positive self-adjoint operator.
-Before presenting the see-saw, let us consider the Cauchy-Schwarz inequality
+where is a Hermitian matrix and the right-hand side of Eq. (72) can immediately be used to define a see-saw algorithm.
+Let us now consider the Cauchy-Schwarz inequality
 where we took into account that . The inequality is saturated for
-Now, the see-saw procedure is as follows:
+Based on these considerations, the see-saw procedure is given as follows.
 (i) Choose some random unit vector.
 (ii) Maximize over and alternatingly. First, is chosen based on Eq. (74). Then, in the alternate step, we use
-We can now recognize the steps of the power iteration, which is known to converge as long as the starting vector ( in our case) overlaps with the eigenvector corresponding to the maximal eigenvalue of .
+With these steps, we realized the power method for finding the largest eigenvalue of a positive self-adjoint operator, which is a classic iterative algorithm in numerical linear algebra [109, 110].
+The iteration is known to converge as long as the starting vector ( in our case) overlaps with the eigenvector corresponding to the maximal eigenvalue of .
+If is not positive semidefinite, then the method converges to the eigenvalue with the largest absolute value.
+Note that the problem could have also been solved with a one shot semidefinite program without iteration.
 V.4 Maximization of the Hilbert-Schmidt norm
@@
 The constraints require that the expectation values of the operators equal the constants The optimization problem will lead to pure states if there are pure states satisfying the condition.
-It can be solved with the method of moments as well as with a see-saw iteration.
+The optimization can be solved with the see-saw method as follows
+where and are density matrices.
 The see-saw method shows a rapid convergence.
 In fact, it converges typically in a single step.
-After examining this simple case, we can have more complicated constraints on the quantum state.
+The problem can also be solved with the method of moments.
+After examining this simple case, in addition to the linear constraints, we can have more complicated constraints on the quantum state.
 For instance, we can consider PPT states or states with a given negativity [10].
@@
 (87).
-The CCNR criterion can detect weakly entangled quantum states, called bound entangle states, that have a positive partial transpose.
+The CCNR criterion can detect weakly entangled quantum states, called bound entangled states, that have a positive partial transpose.
 From this point of view, it is a very relevant question to search for PPT states which violate Eq.
@@
 This can be done by maximizing over PPT states.
-We carried out a numerical optimization, using semidefinite programming.
+We carried out a numerical optimization using semidefinite programming (e.
+g., optimization over PPT states has been considered in Refs.
+[80, 10]).
 We present the maximum of for PPT states for various dimensions in Table 2.
-The state for the case could be determined analytically, see Appendix E.
+The state for the case could be determined analytically, see Appendix F.
 | 1 | |||||
@@
 We discussed how to maximize the metrological gain over local Hamiltonians for a bipartite quantum system.
-We presented methods that exploit the fact that the the quantum Fisher information can be written as a bilinear expression of the elements of the Hamiltonian, if the Hamiltonian is given in the basis of the density matrix eigenstates.
+We presented methods that exploit the fact that the quantum Fisher information can be written as a bilinear expression of the elements of the Hamiltonian, if the Hamiltonian is given in the basis of the density matrix eigenstates.
 We showed an iterative see-saw (ISS) method that can maximize the quantum Fisher information for large systems.
 We showed another method based on quadratic programming.
-We tested the two approaches on concrete examples.
-We also considered several similar problems that can be solved based on such ideas, for instance, obtaining the maximum violation of the Computable Cross Norm-Realignment (CCNR) criterion for states with a positive partial transpose.
+With this so-called relaxation method we can confirm whether the maximum found by the ISS method is indeed a global one for small systems.
+We tested the two approaches on concrete examples, which include one and two copies of the isotropic state of two -dimensional qudits and some important bound entangled states, like, for instance, the Horodecki state.
+These examples also show that we have provided a simple, systematic and efficient method for maximizing the metrological performance (i.e., the metrological gain) of bipartite quantum states with local Hamiltonians.
 Our work can straightforwardly be generalized to multiparticle systems.
+We also considered several similar problems that can be solved based on such ideas, for instance, obtaining the maximum violation of the Computable Cross Norm-Realignment (CCNR) criterion for states with a positive partial transpose, maximizing various norms, searching for the extremal eigenvalues of Hermitian matrices, and also maximizing the Wigner-Yanase skew information.
+Thus, we demonstrated that ISS methods can be used beyond metrological performance optimization in other important quantum information applications.
 VII Acknowledgements
@@
 IT1470-22), and the National Research, Development and Innovation Office NKFIH (Grant No.
-2019-2.1.7-ERA-NET-2021-00036).
+2019-2.1.7-ERA-NET-2021-00036, Advanced Grant No.
+152794).
 We acknowledge the support of the Grant No.
@@
 It is possible to solve Eqs. (93) for and as
-However, if is not known, the recovery is not unique. Eqs. (86) yield a solution for any , for which and holds, constraining as
+However, if is not known, the recovery is not unique. Eqs. (94) yield a solution for any , for which and hold, constraining as
 Note, that the bounds (95) also ensure that . Based on Eq. (95), a necessary condition for the existence of some allowed is
@@
 (97), most of the elements of the coherence vector are zero since
-holds for all Thus, the number of nonzero elements is at most
+holds for all Thus, the number of nonzero elements is at most The remaining elements of the coherence vector are given as
+The following statement can straightforwardly be proven.
+If the state fulfills the condition in Eq.
+(103) then it can be written in the form given in Eq.
+(88).
+Thus, we can decide whether a state is of the form given in Eq.
+(88) based on verifying that Eq.
+(103) holds.
 Appendix B See-saw algorithm with and without semidefinite programming
@@
 Let us try to maximize the convex function
-on a convex set. We could check the extreme point of the set, however, we look for a method that can easily be generalized for larger number of variables.
+on a convex set. We could check the extreme point of the set, however, we look for a method that can easily be generalized to a larger number of variables.
 We introduce another function with an auxiliary variable that is easier to maximize. The linear function tangent to at the point is given as
@@
 We will carry out an optimization of the quantum Fisher information over the convex set of states defined above. Let us now consider the Hamiltonian
-Then, the symmetric logarithmic derivative, defined in Eq. (15), multiplied by is of the form
+Then, we set to the symmetric logarithmic derivative defined in Eq. (15) to
 where is the auxiliary, real parameter. The expression maximized in Eq. (124) equals
@@
 which takes the global maximum for all points for which and Based on these we see that starting from any random initial state, we always reach the global maximum, except for the case when when we obtain However, such states form a zero measure set, thus small additive noise with a random quantum state leads out of the set, and then the algorithm finds the global maximum.
-In the more general case of larger systems, the quantum Fisher information is not the quadratic function of the density matrix elements anymore [111, 112].
+In the more general case of larger systems, the quantum Fisher information is not a quadratic function of the density matrix elements anymore [119, 120].
 However, even in this case, if is maximizing the quantum Fisher information then
@@
 (141).
-Thus, every steady state must fulfill the equation of the given Eq.
-(135) for some Let us consider first the case that has a non-generate spectrum.
+Thus, every steady state must fulfill the condition given in Eq.
+(144) for some Let us consider first the case that has a non-degenerate spectrum.
 Let us look for the eigenvectors of the operator
@@
 This is the case if is not in the range of the initial density matrix.
-However small random added noise can easily help, and the iteration will lead to the global maximum..
+However, small random added noise can easily help, and the iteration will lead to the global maximum..
 Next, we can consider the set of quantum states obtained via a physical map [12].
@@
 (144).
-After a derivation similar to the previous one, we find again that maximizing and the same states maximize it as before.
+After a derivation similar to the previous one, we find again that the algorithm is maximizing and the same states maximize it as before.
 Thus, based on the same arguments, the algorithm always finds the global maximum.
@@
 then the optimization leads to and, based on Eq. (138), we have Thus, it does not reach the maximal value
-The inequalities in Eq. (149) can be added as conditions to an optimization using semidefinite programming. Alternatively, we can also define the a set of quantum states via the physical map, e. g.,
+The inequalities in Eq. (158) can be added as conditions to an optimization using semidefinite programming. Alternatively, we can also define a set of quantum states via the physical map, e. g.,
 where is the projector to the eigenstate of with eigenvalue and is the projector to the eigenstate of with eigenvalue For states coming from this map,
@@
 A family of bipartite PPT entangled states have been found numerically maximizing the quantum Fisher information [10].
-Here, semi-definite programming was used to maximize over the set of PPT quantum states.
+Here, semidefinite programming was used to maximize over the set of PPT quantum states.
 Later, PPT states having the same quantum Fisher information have been found analytically, which is a strong indication that the method found the global maximum [11].
@@
 It would be interesting to prove that the approach always converges for several classes of relevant problems.
-Appendix E Bound entangled state violating the CCNR criterion maximally
+Appendix E An example for using the moment method to upper bound the quantum Fisher information
+In this Appendix, we exemplify the use of the moment method discussed in Sec.
+IV.2 to upper bound the quantum Fisher information for a fixed state .
+We will focus on , however, the method generalizes straightforwardly to any .
+Let us first recap the optimization problem. Fix a state with the eigendecomposition as in Eq. (9). According to Eq. (8), our task is to maximize
+over bipartite local Hamiltonians of the form given in Eq. (24), where we impose the condition
+The matrix elements of in the eigenbasis of are given by .
+Let us denote by the maximum value in Eq.
+(162) over local Hamiltonians.
+Note that the above condition is a relaxation of the stricter one given in Eq.
+(64) for with .
+Therefore, in general defines only an upper bound to the true maximum of the quantum Fisher information.
+The constant coefficients depend only on the eigenvalues of the probe state and are given by Eq.
+(10) whenever the denominator is nonzero.
+Now we apply the following steps to bring the problem to an SDP.
+(i) Build an orthonormal basis of Hermitian matrices. For , we choose the four normalized Pauli matrices , where is the identity, and the rest are , , and .
+(ii) Represent the local Hamiltonian operators as
+with real coefficients , , forming the vector . With this choice we have
+that is, we recover the condition in Eq. (163).
+(iii) For each , define the matrices and , and from them construct the complex vector compactly as
+(iv) In terms of we can write the complex matrix element as
+(v) Hence the quantum Fisher information is
+where the operator
+is a real symmetric matrix, and we used that is real valued.
+This leads us to the optimization problem
+| subject to |
+Thus, we arrived at an optimization task of the form given in Eq. (54). However, we can reformulate this problem as an SDP
+| subject to | ||||
+where we can identify .
+Dropping the non-convex rank constraint yields a convex SDP [see also Eqs.
+(62) and (63)].
+However, it can be shown that the optimum of this relaxed SDP is attained at the rank-one matrix , where is any unit eigenvector of associated with the largest eigenvalue [97, 98].
+Hence the solution to the task (170) is , and the optimal objective value is
+As an illustrative example, let us apply the above procedure to calculate for the two-qubit isotropic state from Eq.
+(61) (see also the first line of Table 1).
+In this case, the matrix from Eq.
+(169) evaluates to
+The eigenvalues of this symmetric matrix are
+Hence, by Eq. (172), we obtain
+One of the eigenvectors corresponding to the maximal eigenvalue can be written as
+where and . Note that since is three-fold degenerate, any unit vector in the corresponding eigenspace gives an optimal .
+From this choice we obtain , and by Eq. (164) the optimal Hamiltonian is given by the bipartite Hamiltonian given in Eq. (24) with
+Incidentally, we find , which also satisfies the original (non-relaxed) constraint of Eq. (64). Thus, in this case the upper bound in (175) is achievable, and the bound is tight.
+Appendix F Bound entangled state violating the CCNR criterion maximally
 In this Appendix, we present the bound entangled state for which the violation of the CCNR criterion given in Eq. (87) is maximal. The state is